# Hard View Selection for Contrastive Learning

### v1

@@ Abstract
Many Contrastive Learning (CL) methods train their models to be invariant to different “views” of an image input for which a good data augmentation pipeline is crucial. While considerable efforts were directed towards improving pre-text tasks, architectures, or robustness (e.g., Siamese networks or teacher-softmax centering), the majority of these methods remain strongly reliant on the random sampling of operations within the image augmentation pipeline, such as the random resized crop or color distortion operation. In this paper, we argue that the role of the view generation and its effect on performance has so far received insufficient attention. To address this, we propose an easy, learning-free, yet powerful Hard View Selection (HVS) strategy designed to extend the random view generation to expose the pretrained model to harder samples during CL training. It encompasses the following iterative steps: 1) randomly sample multiple views and create pairs of two views, 2) run forward passes for each view pair on the currently trained model, 3) adversarially select the pair yielding the worst loss, and 4) run the backward pass with the selected pair. In our empirical analysis we show that under the hood, HVS increases task difficulty by controlling the Intersection over Union of views during pretraining. With only 300-epoch pretraining, HVS is able to closely rival the 800-epoch DINO baseline which remains very favorable even when factoring in the slowdown induced by the additional forwards of HVS. Additionally, HVS consistently achieves accuracy improvements on ImageNet between 0.55% and 1.9% on linear evaluation and similar improvements on transfer tasks across multiple CL methods, such as DINO, SimSiam, and SimCLR.
@@ 1 Introduction
Figure 2: Example images (left) and sampled views (right) after resizing and appearance transformations. Pairs selected adversarially by HVS are shown with a solid frame.
Supervised Learning works that address this shortcoming in a distant manner and which attempt to leverage task difficulty include Spatial Transformer Networks (Jaderberg et al., 2015)and Adversarial AutoAugment (Zhang et al., 2020), which either directly learn geometric transformations or generate augmentation policies that yield hard samples for more robust feature learning. However, these approaches have not been used successfully for unsupervised learning and Adversarial AutoAugment is limited in leveraging the geometric perspective for view hardness control as it does not include RRC in its search space. Furthermore, and in contrast to what we propose, integrating any of these methods into existing contrastive learning pipelines is non-trivial.
Contrastive learning approaches that are related to our work and aim at controlling view hardness include (Shi et al., 2022) and (Tamkin et al., 2021) which propose masking or perturbing input images through adversarial models. Another approach is MADAug (Hou et al., 2023) which use a cooperative setting to learn policies yielding progressively harder samples based on a validation loss which is usually unavailable in contrastive learning. A work that takes a similar perspective like us in leveraging the information content of views for contrastive learning is (Tian et al., 2020b). The authors unify the data augmentation of several CL methods under the lens of mutual information and learn view generators under this premise. While these works offer important insights into view optimality, they usually require learning additional models or non-trivial changes to the pipeline and thus, are not easily integrated into existing CL methods.
In contrast, we propose a fully learning-free, easy-to-integrate approach that applies the following simple strategy: we sample a few views and forward each pair through the currently trained model. Next, we identify the pair that yields the worst loss among all pairs, use it for the backward pass, and repeat (see Fig. 1). While being agnostic to any contrastive loss, this strategy exposes the pretrained model to samples that the model at any given point in the training trajectory finds challenging. As our results show, sampling only four views are sufficient to achieve noticeable improvements.
@@
  • We propose Hard View Selection (HVS), a new learning method complementing contrastive learning that is easy to integrate and extends the broadly-used random view generation to automatically expose the model to harder samples during pretraining. The only requirement of HVS is that it needs to be able to compute sample-wise losses;
  • We demonstrate the effectiveness and compatibility of our approach using ImageNet (Deng et al., 2009) pretraining across popular and widely-adopted contrastive learning methods that cover a diverse range of objectives such as SimSiam (Chen & He, 2021), DINO (Caron et al., 2021), SimCLR (Chen et al., 2020a);
  • HVS improves all baselines noticeably compared to using the common view generation and we show this on a diverse set of transfer datasets and tasks, including linear evaluation, finetuning, object detection and segmentation;
  • We present insights into the underlying mechanisms of HVS. One key observation is that harder samples are achieved by scheduling the Intersection over Union conditioned on the current model state and image input. We also provide further results on the effects of changing the number of sampled views, color augmentation and studies on longer training.
@@ 2 Related Work
2.1 Contrastive Learning in SSL
The core idea behind the contrastive learning framework is to learn image representations by contrasting positive pairs (two views of the same image) against negative pairs (two views of different images) (Hadsell et al., 2006). To work well in practice and to prevent model collapse, contrastive learning methods often require a large number of negative samples (Wu et al., 2018; van den Oord et al., 2018; Chen et al., 2020a; He et al., 2020; Tian et al., 2020a; Chen et al., 2020b) stored in memory banks (Wu et al., 2018; He et al., 2020) or, for instance, in the case of SimCLR, implicitly in large batches (Chen et al., 2020a). Non-contrastive approaches, such as BYOL (Grill et al., 2020), SimSiam (Chen & He, 2021), DINO (Caron et al., 2021) and others (Zbontar et al., 2021; Caron et al., 2020; Ermolov et al., 2021), are able to only use positive pairs without causing model collapse but rely on other techniques, such as Siamese architectures, whitening of embeddings, clustering, maximizing the entropy of the embeddings, momentum encoders and more.
@@ 2.2 Optimizing for Hard Augmentations and Views in SSL
Past work has shown that optimizing augmentation policies or views directly can be beneficial. Hence, the realm of learning task-specific augmentation policies based on data has seen quick development (Cubuk et al., 2019; Ho et al., 2019; Lin et al., 2019; Zhang et al., 2020; Hataya et al., 2020; Müller & Hutter, 2021). Most related to our work among these approaches is Adversarial AutoAugment (AAA) (Zhang et al., 2020) that proposes to generate augmentation policies yielding hard samples by learning with an adversarial supervised objective. In a similar, but cooperative setting, MADAug (Hou et al., 2023) learns policies yielding progressively harder samples based on a validation loss. In contrast to our work, however, AAA and MADAug do not include the random resize crop operation into their search spaces and thus, are limited in leveraging the geometric perspective to control view hardness.
Unlike learning hard augmentations, approaches exist that optimize directly on the pixel level to control view hardness. For instance, Shi et al. (2022) adversarially learn semantically meaningful masks for a pretrained network to inpaint. Another approach perturbs input images using an adversarial model (Tamkin et al., 2021). Tian et al. (2020b) leverage, like us, the view content for better performance. Under the lens of mutual information (MI), they unify several data augmentation CL methods and derive an MI-based objective to learn view generators for unsupervised and semi-supervised learning. While these methods can complement our approach, they are less easily integrated into existing contrastive pipelines due to the need for adversarial network training.
@@ 3.1 Contrastive Learning Framework
In this section, we introduce our approach which is also depicted in Algorithm 12. Many different contrastive learning (He et al., 2020) objectives exist, each characterized by variations stemming from design choices, such as by the use of positive and negative samples or asymmetry in the encoder-projector network structure. For simplicity of exposure, we will introduce our approach based on the SimSiam objective Chen & He (2021), but we do note that our method can be used with any other contrastive learning objective that allows the computation of a sample-wise losses.
@@ 3.2 Hard View Selection
where $\mathcal{L}(\theta)_{i,k,l}$ simply denotes a sample-wise variant of Eq. 1. Overall, we first generate $N$ augmented views for each image $x_{i}$ in the minibatch. Then, we create all combinatorially possible $\binom{N}{2}$ pairs of augmented images for $x_{i}$. Then, we use Eq. 2 to compute the sample-wise loss for each pair which corresponds to invoking forward passes. We then select all pairs that yielded the highest loss to form the new hard minibatch of augmented sets $\textbf{x}^{A_{k_{*}}}$ and $\textbf{x}^{B_{l_{*}}}$, discard the other pairs and use the hard minibatch for optimization. As shown in Algorithm 12, we repeat this process in each training iteration. This modification introduces a more challenging learning scenario in which the model is encouraged to learn more discriminative features by being exposed to harder views. While the additional forward passes come with a certain overhead (about a factor of 1.55$\times$ for SimSiam), as we will demonstrate in our experiments, the focus on harder views leads to consistent improvements.
@@
6:     Create all $\binom{N}{2}$ possible pairs $\textbf{x}_{i}^{A_{k}}\times\textbf{x}_{i}^{A_{l}}$, $k\neq l$
7:     For all pairs $(x_{i}^{A_{k}},x_{i}^{A_{l}})$ forward through $f_{\theta}$ and $g_{\theta}$
@@
While we exemplified the integration of HVS with the SimSiam objective, integrating it into other contrastive methods is as straightforward. The only requirement of HVS is to be able to compute sample-wise losses (in order to select the views with the highest loss). In our experiments section and in addition to SimSiam, we study the application of HVS to the objectives of DINO and SimCLR (see also Appendix H.1 for a formal exemplary definition for the integration of HVS into SimCLR).
@@ 4 Main Results
Here, we discuss our main results on image classification, object detection, and segmentation tasks. All results are self-reproduced using the baseline code and available hyperparameters, and are averaged over multiple seeds. All hyperparameters are reported in Appendix G.
@@ 4.2 Transfer to Other Datasets and Tasks
Table 3: AP50 for object detection and instance segmentation with 100-ep. pretraining.
@@
For object detection, we use the VOC07+12 (Everingham et al., 2010) dataset using Faster R-CNN (Ren et al., 2015). We study instance segmentation on the COCO (Lin et al., 2014) dataset using Mask R-CNN (He et al., 2022). Table 3, where we report the AP50 performance, shows that the features learned with HVS also mostly transfer favourably to these tasks and outperform the SimSiam baseline. However, in case of DINO, we observe a slight performance gap on the detection task, which is the only instance in the entire paper where HVS did not outperform the baseline. This result warrants further investigation, particularly considering that this experiment was conducted with a single seed (due to limited availability of compute resources). More details and performance results on this task are provided in Appendix F.1.
@@ 5.1 Q1: Which Patterns Underlying the Hard View Selection Can be Observed?
When visually studying some of our examples and the views selected by HVS in Figures 2 and 6 (in the appendix), we notice that both geometric and appearance characteristics seem to be exploited. E.g., consider the brightness difference between the views of the first two rows in Fig. 2. Additionally, we observe a generally higher training loss (Fig. 7 in the appendix) indicative of an increased task difficulty. Another key observation is that our adversarial selection chooses view pairs that incorporate zooming in and out or an increased distance between the views (last row in Fig. 2).
@@
To assess these observations quantitatively, we logged relevant hyperparameter data during SimSiam training with HVS and conducted multiple studies which we will describe below. The data we logged include for each view the sampled parameters from the geometric and appearance data augmentation operations (such as the height/width and coordinates of the crops) as well as color distortion parameters (such as brightness and contrast, see Section D in the appendix for more details), as well as the loss and whether the view was selected or not. As evaluated metrics we chose the Intersection over Union (IoU), Relative Distance (image-wise normalized distance of the center points of view pairs), color distortion distance (the Euclidean distance between all four color distortion parameters), and the individual color distortion parameters.
@@
We also examined the IoU value over the course of training in Fig. 4 (left). Here, we notice the average IoU value of selected view pairs increases slightly with training, possibly as a reaction of the pretrained model’s embedding representation to the hard view selection. Another pattern we observe is that the IoU value with HVS (Fig. 4 (left) in blue) is smaller and also less constant during training when compared against training without HVS (green). We believe this is due to the sample-wise and stateful nature of the adversarial selection as HVS chooses different IoU values from sample to sample and training state to training state. Lastly, we assessed the effect of the color augmentation on the pair selection. For this study, we sampled one set of color augmentations (as opposed to multiple, i.e. one each view) per iteration and applied it to all views. Only after identifying the hardest pair, we apply sampled data augmentations for each view individually. Consequently, the linear evaluation performance dropped by 0.3% on average. Interestingly, as we show in Fig. 4 (right), the fraction of selected pairs that are also the hardest pairs slightly increases in this case. One possible explanation for this is that it reflects the non-negligible role of color variation between views (as shown previously with the importance analysis), where HVS is given less leverage to increase hardness through a static appearance and instead, depends more on leveraging the IoU.
@@ 5.2 Q2: Can a Manual Augmentation Policy be Inferred?
Since harder pretraining tasks seem beneficial, and based on the observations made in Q1, a natural question arises: is it possible to mimic the adversarial selection with a manually scripted augmentation policy? Such a policy, if effective, would replace HVS and lower the computational cost by avoiding the additional forward passes. Since the IoU seems to play an important role, below, we study several possible ways for constructing a simple manual augmentation policy based on IoU.
@@
We implemented the following algorithm based on rejection sampling in the augmentation pipeline: we linearly approximate the IoU value curve from Fig. 4 (left; in blue) with a start value of 0.30 and end value 0.35 (ignoring the dip in the early phase) to receive an upper bound for the IoU in each training iteration. For each iteration, we then check if the pairs exceed the IoU upper bound value and if so, we reject the pair and re-sample a new pair (max. 512 times). This ensures that only view pairs are sampled that employ at least a certain task difficulty (by means of a small enough IoU). We varied different hyperparameters (e.g., IoU start/end ranges; inverse schedules; alternating between the IoU schedule and the standard augmentation every other iteration; as well as warmup, i.e. no IoU policy for the first 10 epochs). Training both SimSiam (ResNet-50) and DINO (ViT-S/16; applied to all global and local heads) models for 100 epochs yielded performance decreases or insignificant improvements, as can be seen in Table 4. Moreover, we see an increased risk of model collapse when setting start values too small which can be interpreted as a too hard task. These results indicate that developing a manual policy based on metrics in pixel space such as the IoU is non-trivial. In contrast, we believe that HVS is effective since it 1) operates on a similarity level of latent embeddings that may be decorrelated from the pixel space and 2) has access to the current model state.
@@
To further validate the previous result and to ascertain an upper limit on the performance for predicting the hardest pair, we conducted a second experiment. We fitted a gradient boosting classification tree (Friedman, 2001) to predict the selected view pair conditioned on all SimSiam hyperparameter log data from Q1 (see above) except for the flag that indicates whether a view was selected. As training and test data, we used the logs from two seeds (each 300k samples) and the logs from a third seed, respectively. We also tuned hyperparameters on train/valid splits and applied a 5-fold CV. However, the resulting average test performance in all scenarios never exceeded 40%, indicating that it is indeed challenging to predict the hardness of views based on parameter-level data.
Consequently, we believe these results support the presumption that deriving a policy for controlling and increasing hardness based on geometric and appearance parameters is non-trivial. Furthermore, our analysis suggests that such a policy must function on a per-sample basis and have access to the current model state which is the case when applying HVS.

### v4

# Hard View Selection for Self-Supervised Learning

@@ Abstract
Many Self-Supervised Learning (SSL) methods train their models to be invariant to different “views” of an image and considerable efforts were directed towards improving pre-text tasks, architectures, or robustness. However, most SSL methods remain reliant on the random sampling of operations within the image augmentation pipeline, such as the random resized crop operation. We argue that the role of the view generation and its effect on performance has so far received insufficient attention. To address this, we propose an easy, learning-free, yet powerful Hard View Selection (HVS) strategy designed to extend the random view generation to expose the pretrained model to harder samples during SSL training. It encompasses the following iterative steps: 1) randomly sample multiple views and create pairs of two views, 2) run forward passes for each view pair on the currently trained model, 3) adversarially select the pair yielding the worst loss depending on the current model state, and 4) run the backward pass with the selected pair. As a result, HVS consistently achieves accuracy improvements between 0.91% and 1.93% on ImageNet linear evaluation and similar improvements on transfer tasks across DINO, SimSiam, iBOT and SimCLR. We provide studies to shed light on the inner workings and show that, by naively using smaller resolution images for the selection step, we can significantly reduce the computational overhead while retaining performance. Surprisingly, even when accounting for the computational overhead incurred by HVS, we achieve performance gains between 0.52% and 1.02% and closely rival the 800-epoch DINO pretraining with only 300 epochs.
@@ 1 Introduction
Figure 2: Examples (left) and sampled views (right) after resizing and appearance transformations. Pairs selected by HVS are shown with a solid frame.
Supervised Learning works that address this shortcoming and which attempt to leverage task difficulty include Spatial Transformer Networks (Jaderberg et al., 2015) and Adversarial AutoAugment (Zhang et al., 2020), which either directly learn geometric transformations or generate augmentation policies that yield hard samples. However, these approaches have not been used successfully for unsupervised learning and Adversarial AutoAugment does not include RRC in its search space. In contrast to what we propose, integrating any of these methods into existing self-supervised learning pipelines is non-trivial.
Related SSL approaches exist that aim at controlling view hardness through adversarial Shi et al. (2022); Tamkin et al. (2021) or cooperative Hou et al. (2023) models. A work that takes a similar perspective like us in leveraging the information content of views for SSL is (Tian et al., 2020b). They discover sweet spots for data augmentation hyperparameters that are widely used and adversarially learn flow-based models that generate novel color space-views which are then used for contrastive learning. While these works offer important insights into view optimality, they usually require learning additional models prior to the actual SSL training or non-trivial changes to the pipeline and thus, are not easily integrated into existing SSL methods.
In contrast, we propose a fully learning-free and easy-to-integrate approach. Importantly, and different to Tian et al. (2020b), our approach leverages the current model state and operates on a per-sample basis during pretraining. It applies the following simple strategy: we sample a few views and forward each pair through the currently trained model. Next, we identify the pair that yields the worst loss among all pairs, use it for the backward pass, and repeat (see Fig. 1). While being agnostic to any contrastive and non-contrastive loss, this strategy exposes the pretrained model to samples that the model at any given point in the training trajectory finds challenging. As our results show, sampling only four views are sufficient to achieve noticeable improvements.
@@
  • We propose Hard View Selection (HVS), a new learning method complementing SSL that is easy to integrate and extends the broadly-used random view generation to automatically expose the model to harder samples during pretraining. The only requirement of HVS is that it needs to be able to compute sample-wise losses;
  • We demonstrate the effectiveness and compatibility of our approach using ImageNet (Deng et al., 2009) pretraining across popular and widely-adopted SSL methods that cover a diverse range of objectives such as SimSiam (Chen & He, 2021), DINO (Caron et al., 2021), iBOT (Zhou et al., 2021), and SimCLR (Chen et al., 2020a);
  • HVS improves all baselines noticeably compared to using the common view generation, and we show this on a diverse set of transfer datasets and tasks, including linear evaluation, finetuning, object detection and segmentation;
  • We present insights into the underlying mechanisms of HVS such as the effect of changing the number of sampled views or color augmentation.
@@ 2 Related Work
2.1 Contrastive and Non-Contrastive Learning in SSL
The core idea behind the contrastive learning framework is to learn image representations by contrasting positive pairs (two views of the same image) against negative pairs (two views of different images) (Hadsell et al., 2006). To work well in practice and to prevent model collapse, contrastive learning methods often require a large number of negative samples (Wu et al., 2018; van den Oord et al., 2018; Chen et al., 2020a; He et al., 2020; Tian et al., 2020a; Chen et al., 2020b) stored in memory banks (Wu et al., 2018; He et al., 2020) or, for instance, in the case of SimCLR, implicitly in large batches (Chen et al., 2020a). Non-contrastive approaches, such as BYOL (Grill et al., 2020), SimSiam (Chen & He, 2021), DINO (Caron et al., 2021) and others (Zbontar et al., 2021; Caron et al., 2020; Ermolov et al., 2021), are able to only use positive pairs without causing model collapse but rely on other techniques, such as Siamese architectures, whitening of embeddings, clustering, maximizing the entropy of the embeddings, momentum encoders, and more.
@@ 2.2 Optimizing for Hard Augmentations and Views in SSL
Past work has shown that optimizing augmentation policies or views directly can be beneficial. Hence, the realm of learning task-specific augmentation policies based on data has seen quick development (Cubuk et al., 2019; Ho et al., 2019; Lin et al., 2019; Zhang et al., 2020; Hataya et al., 2020; Müller & Hutter, 2021). Most related to our work among these approaches is Adversarial AutoAugment (AAA) (Zhang et al., 2020) that proposes to generate augmentation policies yielding hard samples by learning with an adversarial supervised objective. In a similar, but cooperative setting, MADAug (Hou et al., 2023) learns policies yielding progressively harder samples. In contrast to our work, however, AAA and MADAug do not include the random resize crop operation into their search spaces and thus, are limited in leveraging the geometric perspective to control view hardness.
Unlike learning hard augmentations, approaches exist that optimize directly on the pixel level to control view hardness. For instance, Shi et al. (2022) adversarially learn semantically meaningful masks for a pretrained network to inpaint. Another approach perturbs input images using an adversarial model (Tamkin et al., 2021). Tian et al. (2020b) leverage, like us, the view content for better performance. Under the lens of mutual information (MI), they derive an objective to adversarially learn flow-based view generators for contrastive learning. While these methods can complement our approach, they are less easily integrated into existing contrastive pipelines due to the need for adversarial network training. Moreover, Tian et al. (2020b) does not use the current model state to control learning hardness and, while their studies incorporate distance metrics, they do not consider the Intersection over Union (IoU) metric which we extensively make use of.
@@ 3.1 Contrastive Learning Framework
In this section, we introduce our approach which is also depicted in Algorithm 12. Many different contrastive learning (He et al., 2020) objectives exist, each characterized by variations stemming from design choices, such as by the use of positive and negative samples or asymmetry in the encoder-projector network structure. For simplicity of exposure, we will introduce our approach based on the SimSiam objective Chen & He (2021), but we do note that our method can be used with any other discriminative SSL objective that allows the computation of a sample-wise losses.
@@ 3.2 Hard View Selection
where $\mathcal{L}(\theta)_{i,k,l}$ simply denotes a sample-wise variant of Eq. 1. Overall, we first generate $N$ augmented views for each image $x_{i}$ in the minibatch. Then, we create all combinatorially possible $\binom{N}{2}$ pairs of augmented images for $x_{i}$. Then, we use Eq. 2 to compute the sample-wise loss for each pair which corresponds to invoking forward passes. We then select all pairs that yielded the highest loss to form the new hard minibatch of augmented sets $\textbf{x}^{A_{k_{*}}}$ and $\textbf{x}^{B_{l_{*}}}$, discard the other pairs and use the hard minibatch for optimization.
@@
6:     Forward all views through $f_{\theta}$ and $g_{\theta}$ to get $z$ and $h$
7:     Create all $\binom{N}{2}$ possible pairs $\textbf{x}_{i}^{A_{k}}\times\textbf{x}_{i}^{A_{l}}$, $k\neq l$
@@
As shown in Algorithm 12, we repeat this process in each training iteration. This modification introduces a more challenging learning scenario in which the model is encouraged to learn more discriminative features by being exposed to harder views. While the additional forward passes come with a certain overhead (about a factor of 1.55$\times$ for SimSiam), as we will demonstrate in our experiments, the focus on harder views leads to consistent improvements.
While we exemplified the integration of HVS with the SimSiam objective, integrating it into other contrastive methods is as straightforward. The only requirement of HVS is to be able to compute sample-wise losses (in order to select the views with the highest loss). In our experiments section and in addition to SimSiam, we study the application of HVS to the objectives of DINO and SimCLR (see also Appendix I.1 for a formal exemplary definition for the integration of HVS into SimCLR).
@@ 4 Main Results
Here, we discuss our main results on image classification, object detection, and segmentation tasks. All results are self-reproduced using the baseline code and available hyperparameters, and are averaged over 3 seeds (iBOT over 2 seeds). All hyperparameters are reported in Appendix G.
@@ 4.2 Transfer to Other Datasets and Tasks
Table 3: AP50 for object detection and instance segmentation.
@@
For object detection, we use the VOC07+12 (Everingham et al., 2010) dataset using Faster R-CNN (Ren et al., 2015). We study instance segmentation on the COCO (Lin et al., 2014) dataset using Mask R-CNN (He et al., 2022). Table 3, where we report the AP50 performance, shows that the features learned with HVS also mostly transfer favourably to these tasks and outperform the SimSiam baseline with a 100-ep. pretraining. However, in case of DINO, we observe a slight performance gap on the detection task, which is the only instance in the entire paper where HVS did not outperform the baseline. This result warrants further investigation, as it was conducted with a single seed (due to limited compute resources). More details and performance results on this task are provided in Appendix F.1.
4.3 Time Complexity of HVS
Method Arch. Lin.
DINO ViT-S/16 74.06
+ HVS ViT-S/16 74.67
Impr. +0.61
SimSiam RN50 68.46
+ HVS RN50 68.98
Impr. +0.52
SimCLR RN50 64.31
+ HVS RN50 65.33
Impr. +1.02
Table 4: Accounting for computational overhead with 50% resolution (100 epochs).
The additional forward passes that HVS introduces for the selection phase increase the time complexity of the individual baseline methods. Several possible approaches exist to mitigate this overhead, one of which is to reduce the image resolution for the selection step (while still using the original resolution for learning). We measured the overhead factors under varying image resolutions for SimSiam and DINO and report them in Table 14 in the appendix. Given the 20% (SimSiam) to 30% (DINO) faster training times per batch when using a 50% image reduction, we evaluated this approach and observed that final performances are retained across all tested methods. Table 4 provides insights into the HVS’ performance when computational overhead is accounted for in the 50% reduction case. Overall, the results validate the effectiveness of HVS in reducing computational demands without compromising on performance. We refer to Section H in the appendix for a discussion on further optimizing HVS.
@@ 5.1 Q1: Which Patterns Underlying the Hard View Selection Can be Observed?
When visually studying examples and the views selected by HVS in Figures 2 and 5 (in the appendix), we notice that both geometric and appearance characteristics seem to be exploited, for instance, see the brightness difference between the views of the first two rows in Fig. 2. We also see a generally higher training loss (Fig. 6 in the appendix) indicative of an increased task difficulty.
@@
To assess these observations, we logged relevant hyperparameter data during SimSiam training with HVS. The logs include for each view the sampled geometric and appearance parameters from the data augmentation operations (such as the height/width of the crops or the brightness; see Section D in the appendix for more details), as well as the loss and whether the view was selected. As evaluated metrics we chose the Intersection over Union (IoU), Relative Distance (normalized distance of the center points views), color distortion distance (the Euclidean distance between all four color distortion parameters), and the individual color distortion parameters.
@@
We also examined the IoU value over the course of training in Fig. 4 (left). An observable pattern is that the IoU value with HVS (Fig. 4 (left) in blue) is smaller and varies more when compared against training without HVS (green). We believe this is due to the sample-wise and stateful nature of the adversarial selection as HVS chooses different IoU values between varying samples and model states. Lastly, we assessed the effect of the color augmentation on the pair selection. For this study, we sampled one set of color augmentations (as opposed to one for each view) per iteration and applied it to all views. Only after identifying the hardest pair, we apply sampled data augmentations to each view. As we show in Fig. 4 (right), the fraction of selected pairs that are also the hardest pairs slightly increases in this case. One possible explanation for this is that it reflects the non-negligible role of color variation between views (as shown previously with the importance analysis), where HVS is given less leverage to increase hardness through a static appearance and instead, depends more on leveraging the IoU. Another key observation is that HVS often chooses view pairs that incorporate zooming in and out or an increased distance between the views (see last row of Fig. 2). We believe the former relates to the analogy of local-to-global view correspondence described in SimCLR and the DINO papers.
@@ 5.2 Q2: Can a Manual Augmentation Policy be Inferred?
Since harder pretraining tasks seem beneficial according to observations made in Q1, a natural question arises: is it possible to mimic the adversarial selection with a manually scripted augmentation policy? Such a policy would replace HVS and lower the computational cost . Since the IoU plays an important role, below, we study several possible ways for constructing a simple manual augmentation policy based on IoU.
@@
We implemented the following rejection sampling algorithm in the augmentation pipeline: we linearly approximate the IoU values from Fig. 4 (left; in blue) with start and end values of 0.30 and 0.35 (ignoring the dip in the early phase). For each iteration, we then check if the pairs exceed the IoU value and if so, we reject the pair and re-sample a new pair. This ensures that only pairs are sampled that entail a minimum task difficulty (by means of a small enough IoU). We varied different hyperparameters, e.g., IoU start/end ranges, inverse schedules, and alternating between the IoU schedule and the standard augmentation . Training both SimSiam and DINO models for 100 epochs yielded performance decreases or insignificant improvements (see Table 5). These results indicate that developing a manual policy based on metrics in pixel space such as the IoU is non-trivial. Additionally, these results show that transferring such a policy from SimSiam to DINO does not work well, possibly due to additional variations in the augmentation pipeline such as multi-crop. In contrast, we believe that HVS is effective and transfers well since it 1) operates on a similarity level of latent embeddings that may be decorrelated from the pixel space and 2) has access to the current model state.
@@
We further validated the previous result and assess an upper limit on the performance for predicting the hardest pair. By fitting a gradient boosting classification tree (Friedman, 2001), we predict the selected view pair conditioned on all SimSiam hyperparameter log data from Q1 except for the flag that indicates whether a view was selected. As training and test data, we used the logs from two seeds (each 300k samples) and the logs from a third seed, respectively. We also tuned hyperparameters on train/valid splits and applied a 5-fold CV. However, the resulting average test performance in all scenarios never exceeded 40%, indicating that it is indeed challenging to predict the hardness of views based on parameter-level data. This further supports our hypothesis that deriving a policy for controlling and increasing hardness based on geometric and appearance parameters is non-trivial and that such a policy must function on a per-sample basis and have access to the current model state as in HVS.